OpenStudy (is3535):

Sixty is what percent of 50? A. 30% B. 90% C. 110% D. 120%

OpenStudy (anonymous):

divide 60 by 50 and multiply by 100%

OpenStudy (is3535):

are you 566666666666666666666666666666666% sure

OpenStudy (anonymous):

60/50*100%=1.2*100%=120%

OpenStudy (anonymous):

yes. I'm sure

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

100% is equal to one

OpenStudy (is3535):

no

OpenStudy (anonymous):

@ksanka is right

OpenStudy (is3535):

it is 120

OpenStudy (anonymous):

so 120% is 1.2

OpenStudy (is3535):

yea bro or girl